If x= t^(1/(t-1)) and y= t^(t/(t-1)) , t>0.t ≠ 1. Then what is the relation between x and y? 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
if we divide Y by X ,
 
then – 
 
y / x = t^[(t –1) / (t –1])
 
y/x = t^1 = t
 
y = tx
 
Regards
Arun (askIITians forum expert)